find the angle equal to 3/7 of its complement.
Answer by nerdybill(7384)

find the angle equal to 3/7 of its complement.
.
Let x = "angle"
then
90-x = complement of angle
.
x = (3/7)(90-x)
multiplying both sides by 7:
7x = (3)(90-x)
7x = 270-3x
10x = 270
x = 27 degrees